UT 20131219 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, MEdwards
Telescope Operator: DGHuerta
LBTO Support:

Plan for the night

- Forecast is not good, probably not much being done tonight
- AGw-transform tests
- LS/MOS-acquisition tests
- Focus-tests (close dome)

Summary

- Closed-dome focus tests

- Accuracy of offsets - tilt of LUCI2-slits at detector differ wrt to the ones
by the telescope when doing detxy offsets (0.7 vs 3deg)

- Throughput of gratings

- Phot ZPs

- Lost 7h due to weather

Preparations

Details

All times are in UT, files are luci2.20131218.iiii

Files taken in the afternoon (no idea who did them) #1-330

00:20 Sunset, overcast at the moment, no wind, hum=10%, T=might improve later on

00:30 Argos team (Gaessler/Peter) did some test for preparation of tomorrow, #331-358

01:50 Focus-tests, closed dome, N1-8-camera at various PA's, K+J
Focus K= -285, J = - 260, always K than J, fwhm @ 960,1020
PA, files, fwhm K, fwhm J 0, #364/365, 5.1/5.04, 4.87/4.96; Out of focus
120, #368/369, 3.1/3.0, 5.0/5.3
180, #370/371, 3.3/3.5, 5.6/5.4
240, #372/373, 3.9/3.4, 6.0/5.4
300, #374/375, 2.9/3.1, 5.2/4.6
360, #376/377, 2.6/2.9, 5.1/4.9
300, #378/379, 3.1/2.9, 5.0/4.7
240, #380/381, 3.9/3.9, 6.1/5.5
180, #382/383, 3.5/3.5, 5.7/5.4
120, skipped
60, #384/385, 3.0/2.9, 5.3/5.0
0, #386/387, 3.0/3.0, 5.2/4.9

Focus K=-325, J=-300
0, #388/389, 1.4/1.0, 1.4/1.6 60, #390/391, 1.4/1.6, 1.4/1.5
150, #392/393, 1.5/1.1, 1.5/1.6
180, #394/395, 1.4/1.4, 1.5/1.6
240, #396/397, 1.2/1.2, 1.9/1.7
300, #398/399, 1.2/1.1, 1.5/1.4
360, #400/401(crap)/402, 1.4/1.4, 1.5/1.4
300, #403/404, 1.5/1.3, 1.4/1.4
240, #405/406, 1.2/1.2, 1.8/1.7
180, #407/408, 1.4/1.2, 1.7/1.6
150, skipped
60, #409/410, 1.4/1.1, 1.5/1.4
0, #411/412, 1.2/1.5, 1.6/1.5

=> Focus should stay on -325

02:23/02:43/02:55/03:16/03:26, Alarm on LUCI2 due to pressure/temp spikes at CCCR1
not clear what the cause is
03:20 Called Walter, reason is insufficient cooling flow...increased, OK

03:01 Error FW2, recovered

03:57 FW2 failing permanently when moving across the K-filter @ rotangle 120
moving to rotangle=210 seems to work, going back to 120 failed immediately,
going to rotangle=300, OK, rotangle=120, fail, rotangle=150, OK,
rotangle=90, fail, rotangle=60, OK

*=> Avoid ROTANGLES=90..140 for moving FW2 across K-filter

05:05 FW2 error at rotangle=360

05:23 ROM hangs when submit a readout

05:52 FW2 error at rotangle 150 (at least twice)

08:30 Weather has improved, cleared up a bit, David checks pointing and collimation
hum went up to 50%, wind to 10 m/s, T=-1deg now, pressure going down (indicating snow tomorrow night)

Accuracy of offsets

Michelle seems to have had problems with the offsets in her Palomar2 field yesterday (smaller than expected),
needs to be checked

08:40 Preset to BS9142, check offsets manually, N3.75, KS, 3*10sec int), #413: pointing (pos=1017,1044=CofR),
#414: offset requested dx=+20", pos=1184,1052, offset found = 20",
#415: offset requested dx=-20", pos=1017,1045, offset found =-20"
#416: offset requested dy=+20", pos=1008,1213, offset found = +20"
#417/418 crap, #419: offset requested dy=-20", pos=1017,1045
=> everything looks ok

Seeing on guider varies between 0.7-1.5", ~0.8" in images

09:10 Preset to BS9142, jittery, mimicking nodding along the slit with 20" offsets in y
in aba(-b)a pattern, repeated 5 times, N3.75, 10*3sec int, #420,421 crap,
#422-429, script stopped after 8th exp., happened now quite a few times during the run
Alexander restarted SW, contn'd, #432/433 with clear/clear, #434-448

AI Alexander/Richard this bug must be removed

Potential problem: tilt of slit is about 0.7deg, while tilt of offset is about 2.7deg
Doesn't really matter for MOS with small offsets but for LS
=> Need to check that

Its clear outside now, lets do the throughput and combine that with check for tilt

Throughput gratings

10:31 Preset to BS9142, #449 sky, #450 src, #451 3"-LS, alignment,
verification, #452, OK, offset dy +40", #453 looks better,
offset dy -80", #454 nada, offset dy +20", #455, a bit of the star,
offset dy +20" #456 (original position), yes, star is there

Confirmed that abs. amount of offset is OK, but that the tilt of
the LUCI slits (0.7deg) and the tilt of the offsets by the telescope (3deg) in detxy wrt to the detector differ
AI LBTO

Seeing degrades, wind gets stronger + humidity rises => storm is approaching

10:58 #457-458 centering star in slit
11:00 #459-462 210z@0.96, #463-466 210J@1.25,#467-470 210H@1.65, #471-474 210K@2.2
11:30 #475-478 200HK@1.93, #479-482 200HK@1.17, #483-486 150Ks@2.17

Seeing >2" now, need to try to correct for substantial flux losses

11:50 Telluric, HIP53735, #487-489 sky, #490-492 src, #493-495 3"-LS, alignment+verification #496-498,
#499/500 210z@0.96, #501/502 210J@1.25, #503/504 210H@1.65, #505/506 210K@2.2
#507/508 200HK@1.93, #509/510 200HK@1.17, #511 150Ks@2.17, than stop (cannot tilt)

Wind close to the limit, seing on guider 2.5-4", about 1.5" on LUCI2-images

12:40 BS9142 in zJHKKs to tie spectra to ZP, N3.75 #512-521 (2 per filter),
N1.8 #522-531 (2 per filter)

13:07 ZPs via BS9126, its photometric but poor seeing, nothing else to do for the last 30mins,
AM 1.75, #532-536 J, #537-541 H, #542-546 K, #547-551 Ks

13:38 Moving to Zenith for flats

#552-561, J, 10*4sec, 4k
#562-571, J, 10*8sec, 12k
#572-581, J, 10*9sec, 20k
#582-591, H, 10*3sec, 7k
#592-601, H, 10*5sec, 12k
#602-611, H, 10*7sec, 24k
#612-621, Ks, 10*3sec, 4k
#622-631, Ks, 10*6sec, 10k
#632-641, Ks, 10*8sec, 15k
#642-651, K, 10*3sec, 9k
#652-661, K, 10*5sec, 15k
#662-671, K, 10*7sec, 25k

14:05 Closing dome

End of COM#2 LUCI2

-- JochenHeidt - 19 Dec 2013
Topic revision: r8 - 17 Jan 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback